# Evaluating Large Multimodal Models for Nutrition Analysis: A Benchmark Enriched with Contextual Metadata

### v1

@@ Abstract
{IEEEkeywords}
Large Multimodal Model, Nutrition Analysis, Portion Estimation, Prompt Engineering.
1 Introduction
[…]
@@ 1 Introduction
2 Related Work
2.1 Traditional Nutrition Analysis Methods
@@
2.2 Image-Based Nutrition Analysis
@@
2.3 Large Multimodal Models for Nutrition Analysis
@@
3 Methodology
3.1 The ACETADA Dataset
@@
3.2 Prompt Engineering for Contextual Metadata
@@
The three contextual metadata flags (gps, timestamp, food) append location information, meal-time information, and specific meal components, respectively. These metadata facets further enrich the meal image with context that the image alone may not convey. The five established prompting modifiers provide a modern baseline of established techniques: classic Chain-of-Thought [50] (cot) elicits step-wise inference; multimodal_cot grounds each numeric guess in the pixels [51]; scale urges explicit size anchoring [52]; few_shot provides numerical exemplars [53]; and expert casts the model as a registered dietitian [8]. Table 1 summarizes each of our available prompting flags.
Table 1: Prompt flags used in our experiments.
@@
4 Results
This section quantifies the impact of contextual metadata and established prompting methods on the accuracy of nutrition and portion-size estimation. For our evaluation, we primarily conduct two experiments. Experiment 1 isolates the stand‑alone benefit of metadata, while Experiment 2 explores how that benefit compounds when metadata is combined with the reasoning-modifying methods introduced in Section 3.2.
4.1 Experimental Setup
@@ 4.2 Model Suite
4.3 Evaluation Protocol and Metrics
[…]
@@ 4.3 Evaluation Protocol and Metrics
Unless otherwise noted, Section 4 shows macro-averaged scores rather than per-model results. Specifically, for any subset of models $M$ (all models, only open-weight, only closed-weight, or any set of prompting methods), the scheme-level MAE is:
@@
4.4 Experiment 1: Impact of Metadata on Simple Nutrition Analysis Prompting
Table 2: Impact of Metadata Integration on LMM Nutritional Estimation Error. This table presents MAE and MAPE changes from a baseline prompt to the best performing metadata combination scheme. Downward arrows ($\downarrow$) and bolded values signify an improvement (i.e., reduced error). The best baseline for each metric is highlighted in blue. Cells are colored green for error reduction and red for error increase, with the largest error reduction being bolded for each nutrient measure.
@@ 4.4 Experiment 1: Impact of Metadata on Simple Nutrition Analysis Prompting
4.5 Experiment 2: Impact of Metadata on Reasoning Modifiers for Nutrition Analysis
To determine whether the contextual block introduced in Section 4.4 can complement more sophisticated prompting strategies, we paired it with five widely used reasoning modifiers: Chain-of-Thought (CoT), Multimodal CoT, Scale-Hint, Few-Shot Exemplars, and Expert-Persona. For each modifier, we compared a baseline prompt, which contained only the meal image and the modifier-specific baseline prompt, with a metadata-enriched prompt where metadata was included from all combinations of gps,timestamp, and food. The most effective combination for every modifier is reported in Table 2.
@@ 4.5 Experiment 2: Impact of Metadata on Reasoning Modifiers for Nutrition Analysis
Table 3: Impact of Metadata Integration on Reasoning Modifier Nutritional Estimation Error. This table presents MAE and MAPE changes from a baseline reasoning modifier to the best performing metadata combination (averaged across all models). Downward arrows ($\downarrow$) and accompanying values signify an improvement (i.e., reduced error). The best baseline for each metric is highlighted in blue. Cells are colored green for error reduction and red for error increase, with the largest error reduction being bolded for each nutrient measure.
@@ 5 Discussion
6 Conclusions and Future Work
@@
7 Supplemental Material
This supplementary document furnishes additional figures and methodological details that extend and contextualize the results reported in the main paper.
7.1 Additional ACETADA Visualizations
To provide a more comprehensive description of the ACETADA dataset, we include two visualizations that are not included in the main text of our study. Figure 9 presents a treemap of the twenty most frequent food items that appear from the dietitian-verified annotations. Figure 10 shows a kernel-density estimate of consumed portion sizes derived from paired before/after meal images. Portions are stratified into “small” ($\leq 500\,$g), “medium” (500500500\,500g-750750750\,750g), and “large” ($>750\,$g) categories. Together, these plots further highlight the diversity of represented foods and the broad range of portion sizes in the ACETADA dataset.
Figure 9: Treemap of the twenty most common food items in ACETADA, as identified by dietitian-verified labels. Block area is proportional to the food item frequency.
Figure 10: Kernel-density estimate of consumed portion sizes in ACETADA, stratified by “small”, “medium”, and “large” consumption categories.
7.2 Experiment 1: Model-Averaged Metadata Effects
Figure 11 illustrates the impact of metadata on MAE and MAPE across all models and metadata combinations. This figure is presented similarly to the radar plots in Figures 6(a)-6(c). With minimal exceptions, incorporating metadata consistently reduces both MAE and MAPE across all models.
(a) Claude-3.7-Sonnet
(b) GPT-4.1
(c) GPT-4o
(d) Gemini-2.5-Pro
(e) Gemma3
(f) Janus-Pro
(g) LLaMA-3.2-Vision-Instruct
(h) Qwen2.5-VL
Figure 11: Averaged Experiment 1 Results For Each Model. MAE (solid lines) and MAPE (dashed lines) are plotted for various contextual metadata combinations for a given model. Each spoke represents an error in a metadata combination; the closer proximity to the center signifies a reduction in error relative to the baseline prompt. Colored markers denote the Best-Metadata configuration for each metric.
7.3 Experiment 2: Modifier and Model-Averaged Metadata Effects
Figures 12 and 13 offer more detailed visualizations of metadata’s effects on nutrient MAE and MAPE across all reasoning modifiers and models. Consistent with the findings in Figure 11, including metadata generally leads to a reduction in nutrient MAE and MAPE. However, it is important to note the exceptionally high MAE values for gps+timestamp and timestamp in the few_shot configuration of Qwen2.5, as depicted in Figure 13(r).
(a) Claude Chain-of-Thought
(b) Claude Expert Persona
(c) Claude Few-Shot
(d) Claude Multimodal CoT
(e) Claude Scale-Hint
(f) GPT-4.1 Chain-of-Thought
(g) GPT-4.1 Expert Persona
(h) GPT-4.1 Few-Shot
(i) GPT-4.1 Multimodal CoT
(j) GPT-4.1 Scale-Hint
(k) GPT 4o Chain-of-Thought
(l) GPT-4o Expert Persona
(m) GPT-4o Few-Shot
(n) GPT-4o Multimodal CoT
(o) GPT-4o Scale-Hint
(p) Gemini-2.5-Pro Chain-of-Thought
(q) Gemini-2.5-Pro Expert Persona
(r) Gemini-2.5-Pro Few-Shot
(s) Gemini-2.5-Pro Multimodal CoT
(t) Gemini-2.5-Pro Scale-Hint
Figure 12: Averaged Experiment 2 Results For Closed-Weight Models. MAE (solid lines) and MAPE (dashed lines) are plotted for various contextual metadata combinations for a given closed-weight model. Each spoke represents an error in a metadata combination; closer proximity to the center signifies a reduction in error relative to the baseline prompt. Colored markers denote the Best-Metadata configuration for each metric.
(a) Gemma3 Chain-of-Thought
(b) Gemma3 Expert Persona
(c) Gemma3 Few-Shot
(d) Gemma3 Multimodal CoT
(e) Gemma3 Scale-Hint
(f) Janus-Pro Chain-of-Thought
(g) Janus-Pro Expert Persona
(h) Janus-Pro Few-Shot
(i) Janus-Pro Multimodal CoT
(j) Janus-Pro Scale-Hint
(k) LLaMA-3.2-Vision-Instruct Chain-of-Thought
(l) LLaMA-3.2-Vision-Instruct Expert Persona
(m) LLaMA-3.2-Vision-Instruct Few-Shot
(n) LLaMA-3.2-Vision-Instruct Multimodal CoT
(o) LLaMA-3.2-Vision-Instruct Scale-Hint
(p) Qwen2.5-VL Chain-of-Thought
(q) Qwen2.5-VL Expert Persona
(r) Qwen2.5-VL Few-Shot
(s) Qwen2.5-VL Multimodal CoT
(t) Qwen2.5-VL Scale-Hint
Figure 13: Averaged Experiment 2 Results For Open-Weight Models. MAE (solid lines) and MAPE (dashed lines) are plotted for various contextual metadata combinations for a given open-weight model. Each spoke represents an error in a metadata combination; the closer proximity to the center signifies a reduction in error relative to the baseline prompt. Colored markers denote the Best-Metadata configuration for each metric.
7.4 Prompt Templates
To facilitate reproducibility, we enumerate the exact text snippets that form our prompts. Each prompt is assembled from three components, presented in order of appearance: (i) a baseline nutrition–analysis request, (ii) an optional reasoning-modifier prefix, and (iii) optional metadata-augmentation fragments. The metadata fragments are appended after the baseline request, whereas the reasoning modifiers are prepended before it.
Table 4 shows the baseline request that is always included. The available reasoning-modifier prefixes are listed in Table 6, and the metadata fragments appear in Table 5. At run time, angle-bracket placeholders (e.g., <Month DD, YYYY>) are replaced with study-specific values before the prompt is sent to the model.
Table 4: Baseline nutrition-analysis prompt (always included).
Prompt (verbatim)
Analyze this food image and estimate the following nutritional content: calories (kcal), protein (g), carbohydrates (g), fat (g), portion (g).
Provide a single point estimate (no ranges) for each metric, using exactly this format:
kcal: <number>, protein: <number>, carbs: <number>, fat: <number>, portion: <number>
Do not include any ranges or hyphens, only one number per metric. ONLY THE NUMERIC ESTIMATES SHOULD BE INCLUDED. THE UNITS FOR EACH METRIC WILL BE ASSUMED; DO NOT INCLUDE THE UNITS.
Table 5: Metadata-augmentation fragments (appended after the baseline prompt). Angle brackets denote run-time substitution.
Flag Prompt fragment (verbatim)
gps Location context: <venue>, <city>, <country>.
timestamp Time context: <Month DD, YYYY at hh:mm AM/PM> (<MealType>, <DayType>).
food Visible foods include: <comma-separated list>.
Table 6: Reasoning-modifier prefixes (prepended before metadata or baseline prompt).
Flag Prompt fragment (verbatim)
expert You are a nutrition expert.
few_shot Analyze the attached meal photo. Here are examples:
Example 1:
 [IMAGE: two slices of pepperoni pizza]
 Output: Calories: ~500 kcal; Protein: ~20 g; Fat: ~22 g; Carbs: ~50 g
Example 2:
 [IMAGE: a bowl of salad with chicken and avocado]
 Output: Calories: ~350 kcal; Protein: ~30 g; Fat: ~15 g; Carbs: ~20 g
[2pt] Now analyze the next image.
multimodal_cot Analyze this meal photo stepwise to avoid mistakes.
1. List the foods present and their approximate amounts.
2. For each item, estimate calories, protein, carbs, fat, and portion size in this exact format (calories-item: X, protein-item: Y, carbs-item: Z, fat-item: W, portion-item: P).
3. Compute and provide the total values in the format described below.
cot Let’s think step by step about the nutritional and portion estimation.
scale Before estimating, please describe what other objects or context you see in the image that could help gauge the size or scale of the food.

### v2

# Comprehensive Evaluation of Large Multimodal Models for Nutrition Analysis: A New Benchmark Enriched with Contextual Metadata

@@ Abstract
I Introduction
Image-based nutrition analysis is increasingly being adopted as a practical and automated alternative to self-report methods such as weighed food records, 24-h recalls (ASA24), and manual logging apps (e.g., MyFitnessPal). Early deep-learning systems—Im2Calories [1], He et al.’s multi-task framework which jointly recognizes foods and infers portion size [2], and the smartphone-centric MUSEFood which fuses RGB-D and inertial cues [3]—demonstrated automatic nutrient estimation with a single meal image, resulting in $\sim\!$15-20 % mean error. Building upon such foundational work, the field is increasingly leveraging Large Multimodal Models (LMMs). These advanced models, which can reason over several different modalities of data, such as images and text, offer substantial promise for further advancing image-based nutrition analysis. However, their application faces two primary challenges:
@@ I Introduction
II Related Work
II-A Traditional Nutrition Analysis Methods
@@
II-B Image-Based Nutrition Analysis
@@
II-C Large Multimodal Models for Nutrition Analysis
@@
III Methodology
III-A The ACETADA Dataset
@@
III-B Prompt Engineering for Contextual Metadata
@@
The three contextual metadata flags (gps, timestamp, food) append location information, meal-time information, and specific meal components, respectively. These metadata facets further enrich the meal image with context that the image alone may not convey. The five established prompting modifiers provide a modern baseline of established techniques: classic Chain-of-Thought [50] (cot) elicits step-wise inference; multimodal_cot grounds each numeric guess in the pixels [51]; scale urges explicit size anchoring [52]; few_shot provides numerical exemplars [53]; and expert casts the model as a registered dietitian [8]. Table I summarizes each of our available prompting flags.
TABLE I: Prompt flags used in our experiments.
@@
IV Results
This section quantifies the impact of contextual metadata and established prompting methods on the accuracy of nutrition and portion-size estimation. For our evaluation, we primarily conduct two experiments. Experiment 1 isolates the stand‑alone benefit of metadata, while Experiment 2 explores how that benefit compounds when metadata is combined with the reasoning-modifying methods introduced in Section III-B.
IV-A Experimental Setup
@@ IV-B Model Suite
IV-C Evaluation Protocol and Metrics
For each nutritional attribute $y\in\{\text{kcal},\text{protein},\text{carbs},\text{fat},$ $\text{portion}\}$ we report the mean-absolute error and the mean-absolute percentage error:
@@ IV-C Evaluation Protocol and Metrics
Unless otherwise noted, Section IV shows macro-averaged scores rather than per-model results. Specifically, for any subset of models $M$ (all models, only open-weight, only closed-weight, or any set of prompting methods), the scheme-level MAE is:
@@
IV-D Experiment 1: Impact of Metadata on Simple Nutrition Analysis Prompting
TABLE II: Impact of Metadata Integration on LMM Nutritional Estimation Error. This table presents MAE and MAPE changes from a baseline prompt to the best performing metadata combination scheme. Downward arrows ($\downarrow$) and bolded values signify an improvement (i.e., reduced error). The best baseline for each metric is highlighted in blue. Cells are colored green for error reduction and red for error increase, with the largest error reduction being bolded for each nutrient measure.
@@ IV-D Experiment 1: Impact of Metadata on Simple Nutrition Analysis Prompting
IV-E Experiment 2: Impact of Metadata on Reasoning Modifiers for Nutrition Analysis
To determine whether the contextual block introduced in Section IV-D can complement more sophisticated prompting strategies, we paired it with five widely used reasoning modifiers: Chain-of-Thought (CoT), Multimodal CoT, Scale-Hint, Few-Shot Exemplars, and Expert-Persona. For each modifier, we compared a baseline prompt, which contained only the meal image and the modifier-specific baseline prompt, with a metadata-enriched prompt where metadata was included from all combinations of gps,timestamp, and food. The most effective combination for every modifier is reported in Table II.
@@ IV-E Experiment 2: Impact of Metadata on Reasoning Modifiers for Nutrition Analysis
TABLE III: Impact of Metadata Integration on Reasoning Modifier Nutritional Estimation Error. This table presents MAE and MAPE changes from a baseline reasoning modifier to the best performing metadata combination (averaged across all models). Downward arrows ($\downarrow$) and accompanying values signify an improvement (i.e., reduced error). The best baseline for each metric is highlighted in blue. Cells are colored green for error reduction and red for error increase, with the largest error reduction being bolded for each nutrient measure.
@@ V Discussion
VI Conclusions and Future Work